OpenStudy (anonymous):

I need to simplifiy using positive exponents. -30x^4y^-4 / 120x^2y^-4

OpenStudy (anonymous):

\[-\frac{x^2}{4}\]

jhonyy9 (jhonyy9):

this is understandably ?

OpenStudy (anonymous):

what do you mean understandably?

jhonyy9 (jhonyy9):

how was simplified ?

OpenStudy (mathteacher1729):

Do you mean? \[\large \frac{-30x^4y^{-4}}{120x^2y^{-4}}\]

OpenStudy (anonymous):

yes, that is how it is set up @ mathteacher

OpenStudy (mathteacher1729):

Ok. Question for you - do you know how to simplify: \(\huge \frac{-30}{120}\) and \(\huge \frac{x^4}{x^4}\) and \(\huge \frac{y^{-4}}{y^{-4}}\) individually?

OpenStudy (mathteacher1729):

sorry, the middle one should be \(\huge \frac{x^4}{x^2}\)

OpenStudy (anonymous):

I can do the -30 and 120 and the x's but the y exponents both being negative throws me off

OpenStudy (mathteacher1729):

So the general rule for dividing exponents is this: \[\huge \frac{x^a}{x^b}=x^{a-b}\] In other words, take the exponent in the numerator and subtract from it the exponent in the denominator. So for \[\huge \frac{x^4}{x^2}=x^{4-2} =x^2\]

OpenStudy (anonymous):

ok, I got that...since -30 goes into 120, I got the answer of -4 and x^2, so far I have the answer of -4x^2 but I am sure of the y's, if they cancel each other out since when you subtract them you get 0 but I know anything to the zero power is =1